OpenStudy (anonymous):

For the function f(x) = x + 4, what is the ordered pair for the point on the graph when x = 3p? (x, x + 4) (x, 3p + 4) (3p, x + 4) (3p, 3p + 4)

OpenStudy (shamil98):

f(3p) = x + 4 f(3p) = 3p + 4 (3p, 3p+4) x = 3p y= 3p + 4